### v1

# Cage: Hardware-Accelerated Safe WebAssembly

Abstract.
WebAssembly (WASM) is an immensely versatile and increasingly popular compilation target. It executes applications written in several languages (e.g., C/C++) with near-native performance in various domains (e.g., mobile, edge, cloud). Despite WASM’s sandboxing feature, which isolates applications from other instances and the host platform, WASM does not inherently provide any memory safety guarantees for applications written in low-level, unsafe languages.
To this end, we propose Cage, a hardware-accelerated toolchain for WASM that supports unmodified applications compiled to WASM and utilizes diverse Arm hardware features aiming to enrich the memory safety properties of WASM. Precisely, Cage leverages Arm’s Memory Tagging Extension (MTE) to (i) provide spatial and temporal memory safety for heap and stack allocations and (ii) improve the performance of WASM’s sandboxing mechanism. Cage further employs Arm’s Pointer Authentication (PAC) to prevent leaked pointers from being reused by other WASM instances, thus enhancing WASM’s security properties.
We implement our system based on 64-bit WASM. We provide a WASM compiler and runtime with support for Arm’s MTE and PAC. On top of that, Cage’s LLVM-based compiler toolchain transforms unmodified applications to provide spatial and temporal memory safety for stack and heap allocations and prevent function pointer reuse. Our evaluation on real hardware shows that Cage incurs minimal runtime () and memory () overheads and can improve the performance of WASM’s sandboxing mechanism, achieving a speedup of over , while offering efficient memory safety guarantees.
1. Introduction
WebAssembly (WASM) (Haas et al., 2017) has been gaining prominence as a versatile compilation target (Musch et al., 2019). WASM allows for deploying and executing native applications written in a variety of languages, such as C/C++ and Rust, in a wide spectrum of environments (e.g., Web, Edge Cloud, IoT devices) (was, 2024b), while achieving near-native performance. In principle, WASM enables the compilation of high-level languages (e.g., C/C++) to its bytecode format, which can then be seamlessly compiled to native machine code based on the targeted underlying architecture, which justifies its increasing adoption.
A core design principle of WASM is the sandboxing of untrusted code (was, 2024a), providing a safety property of Software Fault Isolation (SFI) (Wahbe et al., 1993). Each WASM application is confined in its isolated address space and has no access rights beyond that. Thus, WASM protects the host and other guest WASM instances from potentially malicious or buggy code.
However, applications compiled to WASM are still vulnerable to memory safety issues, such as buffer overflows or dangling pointers, within an application’s memory space, despite WASM’s sandboxing (Lehmann et al., 2020). Such issues allow attackers to manipulate the memory space of a WASM instance, corrupting or leaking sensitive data or manipulating the control flow. This limitation becomes particularly evident when compiling applications written in memory-unsafe languages, e.g., C/C++. Importantly, numerous CVEs remain exploitable when compiled to WASM (3).
Unfortunately, existing memory safety solutions, designed for applications compiled to native machine code, cannot be directly applied for WASM due to several factors:
Memory safe languages, such as Rust or OCaml, tackle memory safety at the language level by encoding certain properties in their type system or providing safe abstractions, e.g., in their standard libraries. However, it is not feasible to port all applications or libraries written in C/C++ to languages with such guarantees.
Pointer- or object-based solutions (Nagarakatte et al., 2009; Kroes et al., 2018; Kuvaiskii et al., 2017; Akritidis et al., 2009; Nethercote and Seward, 2007; Michael et al., 2023; Duck and Yap, 2016; Dhurjati et al., 2006; Duck et al., 2017) can be adapted to WASM, but they either modify the pointer layout and size, incur significant runtime overheads, or rely on custom hardware (Watson et al., 2020), thus, hindering production deployment in common WASM environments.
Hence, we aim to answer the question: How can we design a practical system to provide memory safety properties for unmodified applications compiled to WASM without changing its linear memory model and introducing minimal overheads to allow for deployment in production environments?
Our key insight is to leverage modern commodity hardware extensions by designing abstractions for WASM that can be directly used by the compilers. Precisely, Arm recently introduced new ISA extensions, namely PAC (Qualcomm Technologies, Inc., 2017) and MTE (ARM Ltd., 2019), that can serve as a base to design practical, high-performance solutions against memory safety issues.
To this end, we propose Cage, a hardware-accelerated WASM toolchain that leverages Arm’s MTE and PAC hardware extensions to provide both spatial and temporal memory safety issues for unmodified C/C++ programs. Cage further hardens applications against control flow highjacking in the form of function pointer reuse between WASM instances. Additionally, Cage improves the performance of the sandboxing mechanism for 64-bit WASM by offloading the bounds checks to MTE hardware. We design Cage so it can fall back to software-based implementations on devices lacking the relevant hardware.
To the best of our knowledge, Cage is the first practical, comprehensive, and efficient solution for unmodified C/C++ programs that can be deployed on commodity hardware and provides strong memory safety guarantees for WASM instances running memory-unsafe code on Arm platforms.
Altogether, Cage makes the following contributions:
WebAssembly Extension: A minimal and generic extension to the WebAssembly specification to provide memory safety guarantees and is deployable on every platform, regardless of the availability of specialized hardware.
Compiler Toolchain: A compiler toolchain that transparently hardens unmodified programs to enforce spatial and temporal memory safety for stack and heap allocations and prevent unintended function pointer reuse.
WASM Runtime: A hardware-accelerated WASM compiler and runtime, leveraging Arm’s MTE and PAC with minimal overhead that can be deployed in production.
Evaluation on real hardware: Evaluation and security analysis of Cage’s implementation on commercially available Arm hardware. Our evaluation is structured around performance and memory overheads and is accompanied by an extensive analysis of MTE and PAC performance as implemented on real hardware.
We implement our Cage prototype on top of 64-bit WASM, consisting of a compiler toolchain based on LLVM and a WASM compiler and runtime based on wasmtime incorporating support for Arm’s MTE and PAC. We conduct an extensive analysis of Cage’s security guarantees and evaluate Cage’s performance using the PolyBench/C (Pouchet, [n. d.]) benchmark suite. We further analyze Arm’s MTE and PAC performance as implemented on production hardware through a set of microbenchmarks. Our evaluation on real Arm hardware shows that Cage provides its memory safety properties while incurring minimal runtime () and memory () overheads and is capable of significantly improving the performance of WASM’s sandboxing feature, achieving a speedup of over .
2. Background
2.1. WebAssembly
WASM’s linear memory model.
[…]
WASM also provides sandboxing for programs. The WASM runtime must ensure that each instance can only access memory within the bounds of its accessible linear memory. This is typically achieved using either (i) explicit bounds checks that are inserted before every memory access, validating that it lies within the WASM instance’s memory range or (ii) guard pages, where 4 GiB of virtual memory is mapped and pages beyond the WASM instance’s memory is marked as inaccessible, and any access in them results in a segmentation fault. This only works for 32 bit pointers, which don’t allow accessing memory beyond 4 GiB. Typically, switching to 64-bit WASM entails switching to the more expensive approach (i) with explicit bounds checks.
, while the surrounding memory is tagged with
. When accessing memory, the hardware catches out-of-bounds errors (). When freeing memory, the memory region is retagged with
. This prevents use-after-free errors ().2.2. Memory Safety
Applications written in low-level, memory-unsafe languages (e.g., C/C++) are prone to memory safety bugs that enable a whole class of attacks on a vulnerable or buggy program (Szekeres et al., 2013). Several studies have shown that in large software projects, memory safety bugs make up between 70 % and 75 % of their security vulnerabilities (chr, [n. d.]; Thomas, [n. d.]; Vander Stoep and Zhang, [n. d.]). These bugs are classified as spatial or temporal. Spatial memory safety errors occur when a memory access is performed beyond the allocated boundaries of a memory object (e.g., buffer overflows), while temporal memory safety bugs refer to accesses to memory regions before they are allocated or after their release (e.g., dangling pointers). These vulnerabilities can be exploited by attackers to overwrite data or redirect control flow, for instance, by manipulating the return address saved on the stack to create return-oriented programming (ROP) chains, chaining existing snippets of code together to create attacks.
To this end, several memory safety solutions have been proposed, which can be divided into three major categories: (i) trip-wire-based approaches (Serebryany et al., 2012, 2023; Bozdoğan et al., 2022; Nethercote and Seward, 2007; Hasabnis et al., 2012; Dang et al., 2015) that employ guard zones around memory objects and allocate designated memory regions, namely shadow memory, that determine whether a memory location is accessible or not, (ii) object-based approaches (Akritidis et al., 2009; Duck and Yap, 2016; Dhurjati et al., 2006; Duck et al., 2017) that track memory safety metadata on a per-object level and ensure memory safety for pointers with respect to an object, and (iii) pointer-based approaches (Nagarakatte et al., 2009; Necula et al., 2002; Jim et al., 2002; Kroes et al., 2018; Corporation, 2013; Woodruff et al., 2014; Kuvaiskii et al., 2017) that keep track of object bounds by either storing them in the pointer itself (e.g., via fat pointers) or in external data structures. Additionally, specialized solutions (Abadi et al., 2009; Kuznetzov et al., 2018; Liljestrand et al., 2019) have also been proposed to mitigate control-flow attacks that occur via memory safety bugs and either analyze applications to enforce valid control flow paths or enforce memory safety for code pointers using hardware- or software-based techniques.
Memory tagging approaches (Serebryany et al., 2018; ARM Ltd., 2019) combine aspects from object- and pointer-based approaches. They typically associate metadata stored in the unused bits of a pointer with memory objects by assigning tags to allocated memory regions and performing checks at runtime.
Pointer layout on aarch64 in Linux with and without MTE and PAC enabled.
2.3. Memory Safety Hardware Extensions
As an alternative to flexible yet slow software-based memory safety solutions, CPU designers develop hardware extensions (ARM Ltd., 2019; Qualcomm Technologies, Inc., 2017; Corporation, 2013) to serve as a foundation for efficient memory safety. They provide security primitives to fortify applications with memory safety properties while having minimal memory and performance overheads, making them suitable for production deployment.
Fig. 3 presents an example layout of a pointer on aarch64, the 64-bit variant of the ArmV8 ISA (ARM Ltd., [n. d.]). Only 48 out of the available 64 bits are used to address memory. The remaining bits are set to 0 or 1 to differentiate between kernel and userspace. Hardware extensions such as Top Byte Ignore (TBI), MTE (2.3), or PAC (1) leverage those unused bits for storing metadata.
[…]
MTE currently can be set in one the following modes: (i) disabled, where no tag checks are performed, (ii) synchronous, where a tag mismatch immediately triggers a fault disallowing the read/write at the affected memory location, (iii) asynchronous, where a tag mismatch does not cause a fault directly but sets a CPU flag that is checked at the next context switch, thus allowing for a potential read/write at the affected memory location by the triggering command, and (iv) asymmetric, where reads are checked asynchronously and writes are checked synchronously.
| Inst | Cortex-X3 | Cortex-A715 | Cortex-A510 | |||
| Tp | Lat | Tp | Lat | Tp | Lat | |
| MTE | ||||||
| irg | 1.34 | 1.99 | 1.00 | 2.00 | 0.50 | 3.00 |
| addg | 2.01 | 1.99 | 3.81 | 1.00 | 2.22 | 2.00 |
| subg | 2.01 | 1.99 | 3.81 | 1.00 | 2.22 | 2.00 |
| subp | 3.49 | 0.99 | 3.81 | 1.00 | 2.50 | 2.00 |
| subps | 2.88 | 0.99 | 3.80 | 1.00 | 2.50 | 2.00 |
| stg | 1.00 | – | 1.81 | – | 1.00 | – |
| st2g | 1.00 | – | 1.84 | – | 0.46 | – |
| stzg | 1.00 | – | 1.84 | – | 0.98 | – |
| st2zg | 0.34 | – | 1.79 | – | 0.45 | – |
| stgp | 1.00 | – | 1.69 | – | 0.98 | – |
| ldg | 2.92 | – | 1.91 | – | 0.93 | – |
| PAC | ||||||
| pacdza | 1.01 | 4.97 | 1.51 | 5.00 | 0.20 | 4.99 |
| pacda | 1.01 | 4.97 | 1.42 | 5.00 | 0.20 | 5.00 |
| autdza | 1.01 | 4.97 | 1.51 | 5.00 | 0.20 | 7.99 |
| autda | 1.01 | 4.97 | 1.43 | 5.00 | 0.20 | 7.99 |
| xpacd | 1.01 | 1.99 | 1.56 | 2.00 | 0.20 | 4.99 |
[…]
MTE and PAC can be combined at the cost of bits available for the PAC signature. The exact layout of the PAC signature varies depending on the system. On Linux, bits 56–59 are used for MTE while bits 63–60 and 54–49 are used for PAC, as shown in Fig. 3. The remaining bit 55 differentiates between kernel- and user-space addresses.
Architectural performance analysis. We evaluate the performance characteristics of MTE as implemented on the Tensor G3 chip, such as throughput and latency of the individual instructions in Table 1. We run microbenchmarks executing instructions in an unrolled loop. To measure throughput, we execute the instructions without any data dependencies; to measure latency, we force a data dependency between subsequent instructions. For instructions storing and loading memory tags, we only measure throughput.
We further measure the raw overhead of enabling MTE. We perform a 128 MiB memset with MTE disabled, synchronous, and asynchronous mode. We perform each run with a clean cache. In Fig. 4, we observe that with synchronous MTE, memset is 19.1%, 14.4%, and 29.9% slower on the respective cores compared to the baseline with MTE disabled. Asynchronous MTE gets closer to the baseline with an overhead of 2.6%, 3.3%, and 11.3%, respectively.
3. Motivation
WebAssembly is inherently protected by design against a wide range of attack vectors, such as jumping to arbitrary addresses or injecting shellcode. Despite its protection mechanisms, it is shown that WebAssembly is still susceptible to attacks originating from memory safety issues, such as buffer overflows or dangling pointer accesses (Lehmann et al., 2020).
Importantly, WASM compilers place data in the linear memory of the WASM instance, where both read and write permissions are granted to the executed code. The lack of read-only memory regions and the ability to map arbitrary pages in WASM prevents measures, such as Address Space Layout Randomization (ASLR) or guard pages, from being applied. Thus, in case of a successful exploit of a memory safety bug, an attacker can overwrite application data. Table 2 presents a set of exploits that were discovered outside the scope of WASM but serve as examples that showcase potential memory safety errors when the vulnerable applications run in a WASM instance.
| CVE | Cause | Mitigated in WASM |
| CVE-2023-4863 | Out-of-bounds | No |
| CVE-2014-0160 | Out-of-bounds | No |
| CVE-2021-3999 | Out-of-bounds | Partially11footnotemark: 1 |
| CVE-2018-14550 | Out-of-bounds | No |
| CVE-2021-22940 | Use-after-free | No |
| CVE-2021-33574 | Use-after-free | No |
| CVE-2020-1752 | Use-after-free | No |
| CVE-2019-11932 | Double-free | Partially††footnotemark: |
Additionally, in WASM, a degree of control flow manipulation is possible. Precisely, function pointers can be overwritten with pointers to other functions that share the same signature. However, it can only occur with functions that are present in the function table, i.e., functions that are targets of virtual function calls, such as virtual functions in C++. This results in a similar property as code-pointer separation (CPS) (Kuznetzov et al., 2018). LABEL:lst:vulnerable-overflow-pac showcases such a vulnerability, as an attacker can overwrite a function pointer and redirect an indirect call to another function. When WASM is used in a WebOS-like scenario (kwast os, 2024; Krieger, 2024; Wen and Weber, 2020), i.e., running multiple instances in a single process, leaking function pointers in one program is even more vulnerable, especially if instances share a common library.
[…]
WASM is, despite limitations such as high overhead when running 64-bit WASM or the lack of a practical and low-overhead solution to prevent memory safety issues in programs written in C/C++, steadily growing in adoption (was, 2024b).
Recently, manufacturers have started shipping hardware supporting both PAC and MTE, with the Google Pixel 8 being the first commercially available device to feature both.
These extensions offer strong security guarantees with very low overhead, as shown in Table 1 and Fig. 4.
4. Design
Cage consists of a WASM extension to provide memory safety guarantees within a WASM instance and improve the performance of the sandboxing mechanism. Its core design principles are to be (i) minimally invasive and (ii) applicable on diverse platforms using various approaches, including hardware extensions, such as MTE and PAC, software-based techniques, or hybrid solutions similar to HWASan (Serebryany et al., 2018).
Figure 5 presents an overview of Cage. Precisely, unmodified C/C++ source code is compiled using the LLVM toolchain (Lattner and Adve, 2004), where Cage includes a sanitizer that instruments stack allocations and function pointers, along with a modified libc based on WASI (Group, 2024b) that protects heap allocations. LLVM’s backend generates the hardened WASM binaries that can be executed in wasmtime (Alliance, 2024b), which implements our extension using MTE and PAC.
4.1. System Model
Threat model.
Figure 6 highlights two aspects of memory safety, present in WASM, namely internal and external memory safety.
We depict trusted components (
) in green and untrusted components (
) in red.
We further annotate the hardened component by Cage in each of these models (
The internal memory safety model (Fig. 6(a)) mirrors the threat environment of a standard non-WASM program. In this case, the application in the sandbox and its runtime, including the compiler, are considered trusted and assumed to be bug-free. Untrusted input (e.g., network data) originates outside the sandbox and can be controlled by an attacker. This implies that common memory safety bugs, such as buffer overflows or use-after-free, can be exploited to tamper with the WASM memory. WebAssembly’s design inherently mitigates some threats common in non-WASM environments ( 2.1), so we do not consider ROP-style attacks or those relying on unstructured control flow.
The external memory safety model (Fig. 6(b)) refers to the security of the sandbox. Threats originate from running untrusted programs, which may be adversarial or contain bugs. Typical attacks include sandbox escapes, where an attacker attempts to break out of the sandbox’s restrictions and access host resources, or side-channel attacks, where attackers exploit timing differences or resource usage patterns to infer sensitive information. In this setting, we assume that the operating system and the underlying architecture do not have exploitable bugs. Importantly, we do not make assumptions about potential spectre-like (Kocher et al., 2020) attacks. We do not protect against protected against side-channel attacks.
Programming model. Cage currently supports unmodified applications written in C/C++ that target 64-bit WASM. Note that our toolchain is not bound to any language and can handle other languages compiled to LLVM in the future. Cage provides spatial and temporal heap safety to applications that use its adapted WASI-libc that comes with a hardened allocator. For applications using their own allocator, we expose Cage’s memory safety primitives to C, enabling programmers to implement the same security guarantees. Lastly, Cage reserves the unused upper 16 bits of 64-bit pointers to place memory safety metadata.
Deployment model. Cage’s prototype is highly optimized to use Arm’s hardware extensions to minimize the runtime overheads. Therefore, its primary deployment target is Arm CPUs with Arm PAC and MTE extensions. However, Cage can also be deployed on any platform, regardless of the underlying hardware, where the respective memory safety protection mechanisms have an equivalent software fallback, resulting in higher performance overheads.
| i64.pointer_sign | |||
| i64.pointer_authenticate |
4.2. WebAssembly Extension
Cage essentially is an extension to WebAssembly that introduces primitives that can be used by a modified standard library or sanitizers to provide memory safety guarantees for selected memory allocations. It builds on wasm64 (Group, 2024a), the 64-bit variant of WebAssembly. Wasm64 extends pointers within a WASM instance to 64 bits. Out of those, only 48 are used to index memory. This allows Cage to utilize the remaining 16 bits to store its required memory safety metadata.
Memory safety. Cage introduces the notion of segments and tagged pointers in the context of WASM. It further features five new instructions, shown in Fig. 7, that allow the creation of segments and the derivation of tagged pointers from raw pointers. The Cage pointers carry provenance and can only access the segment they were created with. Conversely, segments can only be accessed by the tagged pointers created with them.
At an instance startup, the linear memory consists of a single segment that can be accessed via untagged indices, allowing unmodified code to run under our new semantics without modifications. This design choice allows the gradual integration of safety primitives into specific parts of WebAssembly applications where enhanced security is required. Thus, Cage can achieve security properties similar to those of specialized hardware designed mainly for memory safety (Woodruff et al., 2014) by combining the concept of segments with Arm’s hardware extensions (e.g., MTE).
Heap safety. In Cage, the memory allocator must be aware of segments to provide heap safety. Adapting a memory allocator to utilize Cage’s memory safety extensions requires minimal modifications. When allocating memory, the allocator must align the requested size to 16 bytes, perform the allocation, and create a segment. The corresponding tagged pointer is then returned to the caller. Cage randomly selects a tag for each allocation. When freeing memory, Cage’s allocator uses the segment.free instruction that ensures the detection of potential use-after-free and double-free errors.
Further, Cage has to ensure that there are no tag collisions for adjacent allocations to provide protection against off-by-one buffer overflows/underflows. This is achieved by design as Cage places metadata at the beginning of every heap allocation, and the corresponding memory region is preserved untagged, as shown in Fig. 8(a). Thus, adjacent allocations are always separated by an untagged memory segment.
Stack safety. To provide memory safety for the stack, Cage creates segments for the stack slots when entering a function. Cage generates a random tag per function for the first stack allocation. Subsequent stack allocations use this tag and increment it by one. As the available tag bits are limited, the tag wraps around on overflow. Before returning from a function, all stack slots are untagged and reassigned to the original stack frame. This allows other functions to use the memory and prevents stack slots from being accessed after returning from a function. Note that each stack allocation needs to be aligned to 16 bytes and gets processed when entering/exiting from a function.
Similarly to the heap, Cage must guarantee protection for off-by-one overflows/underflows on the stack. Therefore, Cage must ensure that two adjacent stack allocations between functions do not share the same tag. To achieve this, Cage inserts a single untagged stack guard slot at the beginning of the frame if no such untagged stack slot exists, as shown in Fig. 8(b). Without the guard allocation, adjacent allocations in stack frames and would share the same tag (blue) and, thus, Cage would not be able to detect overflows.
To further reduce its performance and memory overheads, Cage omits the instrumentation of stack allocations that (i) do not escape the function or (ii) are only accessed using statically verifiable indices. In Algorithm 1, we present a simplified version of Cage’s algorithm for the identification of the (un)safe stack allocations.
Pointer authentication.
Cage provides pointer authentication primitives that prevent function pointer reuse between WASM instances.
On the instantiation of a WASM module, a secret key is generated. The key is not accessible by the user code.
Cage’s authentication operations leverage this key to sign and authenticate pointers using a cryptographic hash function.
The signature is placed in the unused 16 bits of a WASM pointer, alongside the pointer tag, if applicable.
Pointers containing a signature cannot directly access memory.
On authentication, the signature is checked and stripped, if it is valid. Otherwise, the WASM module traps.
While function pointer reuse within a WASM instance is still possible, Cage prevents the reuse across different instances, as each instance generates its own key.
While memory64 (Group, 2024a) extends function pointers to 64 bits, the indices for the WASM function table remain 32 bit wide.
Cage uses the instruction sequence in Fig. 9 to authenticate function pointers and perform indirect function calls.
64-bit pointers are first authenticated, which traps in case of a invalid signature.
If successful, the signature is stripped and the pointer is truncated to 32 bits.
Similarly, when creating function pointers, indices into the function table are first zero-extended to 64 bits and then signed.
5. Semantics
5.1. Typing Rules
Cage extends the typing rules of the original WASM paper (Haas et al., 2017), as shown in Figure 10. We adopt the notation used in previous work (Pierce, 2002). Specifically, the rules are of the form of . An instruction is valid under the context , with being used to access a context component, such as the memory. The rule ensures that the instruction can only be used when memory is declared. The type describes how the instruction manipulates the operand stack. The instruction expects an operand stack where it pops and pushes .
| (2) | otherwise | ||||
| (4) | otherwise | ||||
| (6) | otherwise | ||||
| (8) | otherwise | ||||
| (10) | otherwise | ||||
| (13) | otherwise | ||||
5.2. Small-step Reduction Rules
Figure 11 highlights how Cage extends the WASM small-step reduction rules (Haas et al., 2017) using the notation established by previous work (Plotkin, 1981). The lower part of Fig. 11 presents the new tag-aware load/store rules that replace the load/store rules from the WASM paper as well as new rules tailored based on the introduced instructions. To signal a trap, we reuse operators from the original WASM rules, including the trap operator. The store, , is augmented with a storage mechanism that assigns a tag, , to each 16-byte memory granule and a per-instance secret key . The key is unique per WASM instance and ensures that leaked signatures cannot be used in another instance or in another run of the same instance. Precisely, we use the following notation:
: Extracts the tag for a memory region in instance accessed at address with length , if the tag is the same for all bytes in the range .
: Updates the state with new tags for the memory region at address with length , if is aligned to 16 bytes and the memory region is in bounds of the memory.
: Extracts the tag from a tagged pointer.
: Creates a tagged pointer from an untagged pointer to be used for a new segment. The tag is randomly chosen from a pool of tags.
: Creates a tagged pointer for the purpose of freeing a segment. The tag is different from the tag stored in .
: Creates a cryptographic signature based on the index and a per-instance secret key and inserts it into the upper bits of .
: Removes the cryptographic signature from the upper bits of .
Further, Figure 11 presents the added and modified components in the rules. Each reduction rule is depicted with the top of the operand stack and the state, , on the left side, representing the pre-execution state. The resulting stack and state after the execution of the instruction is placed on the right side. For rules with , represents the instance in which the instruction is executed. Cage modifies the semantics of load/store instructions by adding new rules for loads and stores in Eqs. 1, 2, 3 and 4, which trap on tag mismatches.
Additionally, Cage introduces new rules that express its new instructions, which are described below:
segment.new: It creates a new, zeroed memory segment for a given pointer and size and returns a tagged pointer (Eq. 5).
segment.set_tag: It transfers ownership for a given pointer and size to a tagged pointer (Eq. 7). This results in the tagged pointer being able to access the memory segment and can be used to merge adjacent segments.
segment.free: It invalidates a segment specified by a pointer and length by tagging the segment with a new, implementation-defined tag (Eqs. 9 and 10), ensuring accesses to the freed segment are caught. This instruction also traps if a segment is freed twice, i.e., the given tagged pointer cannot access the memory region.
i64.pointer_sign: It signs a pointer and places a cryptographic hash in its upper bits (Eq. 11).
Using the above rules, the Eqs. 6, 8 and 10 of Figure 11 ensure traps when code tries to create, modify, or free unaligned segments or segments outside the linear memory for a given instance. Cage requires all segments to be aligned to 16 bytes.
6. Implementation
Cage is built based on LLVM 17, wasi-libc, and wasmtime 16.0.0. Cage consists of a modified LLVM toolchain (6.1), wasi-libc (6.2) and implementation of the memory safety extension in wasmtime (6.3, 6.4).
6.1. LLVM Compiler Toolchain
In Cage, we use LLVM (Lattner and Adve, 2004) to compile our C/C++ applications to WebAssembly. We extend its existing WASM backend to be able to emit the new Cage’s instructions. We further introduce new intrinsic functions that correspond and are lowered to Cage’s WASM instructions by the compiler. The clang frontend and the Cage compiler passes are mainly responsible for inserting calls to Cage’s new intrinsic functions when necessary. Precisely, we introduce two WASM-specific sanitizer passes that can be enabled in the LLVM pipeline via compiler flags. In clang, we introduce new built-in functions, which directly map to these intrinsics and allow programmers to use the new instructions, e.g., to build a segment-aware memory allocator.
The first sanitizer pass is designed to provide memory safety for stack allocations when compiling the source code to WebAssembly. It analyzes functions for stack allocations, applies padding, and creates the memory segments (8), ensuring temporal and spatial safety of stack allocations.
The second sanitizer pass enforces pointer authentication for indirect function calls. Cage instruments code taking references to functions and performing indirect calls. We do not handle other operations on code pointers, as pointer arithmetic on function pointers is undefined behavior in C (International Organization for Standardization, 2018, §6.5.6/2, §6.5.2/1) and C++ (International Organization for Standardization, 2020, §7.6.6/1, §6.8/8, §31.8.4/4).
Note that both sanitizer passes run after all LLVM optimizations. This ensures that Cage does not block passes that might remove stack allocations, such as mem2reg.
6.2. WASI Libc Modifications
We port the WebAssembly System Interface (WASI) and wasi-libc to wasm64 to run applications relying on libc on wasm64. To achieve this, Cage adapts the size and the pointer types from 32 to 64 bits. Further, we modify dlmalloc, the default allocator in wasi-libc, to provide memory safety for heap allocations. Cage’ allocator creates memory segments and returns tagged pointers to these segments. Thus, it protects both allocator metadata and adjacent allocations from being accessed or modified through heap overflows, as illustrated in Fig. 8(a). When freeing or reallocating memory, segments are freed, ensuring temporal safety. Lastly, we recompile WASI-libc with pointer authentication to ensure function pointers are signed and authenticated by library code.
6.3. Internal Memory Safety
Figure 12(a) illustrates an overview of Cage’s implementation of the internal memory safety (4.1) extension in wasmtime using MTE. We modify wasmtime and its supporting libraries to parse, process, and enforce the memory safety extension described in 4.2. More specifically, we add support for MTE and PAC in the form of new instructions and lowering rules. We implement segments using MTE. Cage provides memory safety for the heap via its modified allocator that protects heap allocations by creating segments. Respectively, to provide memory safety for the stack, Cage adapts the stack allocations; the first segment in each function is assigned a random tag and successive allocations increment the tag by one (4.2), eventually wrapping around. This ensures adjacent allocations on the stack never share a tag. Importantly, Cage uses synchronous MTE to trap on memory safety violations before their effects become observable.
For pointer authentication, Cage ensures that each WASM instance receives a distinct secret key to sign pointers. If multiple WASM instances run in a single process, Cage initializes a global random value per instance used as the PAC’s instructions modifier. This is required, as PAC keys are shared per process. Otherwise, Cage uses PAC instructions without a modifier.
6.4. External Memory Safety
Fig. 12(b) presents how Cage utilizes memory tagging to replace software-based bounds checks and preserve external memory safety (4.1) for each WASM instance. Initially, the runtime assigns a tag to each instance on module instantiation, which is stored in the heap base address. The effective memory access address is calculated by adding the accessed index to the tagged heap base address. The memory of the runtime is tagged with zero. It enables MTE to catch any access outside the sandbox due to the tag mismatch.
Cage can further combine both its external and its internal memory safety extensions by splitting the pointer tag bits among them. The upper bits are used for internal memory safety, while the lower bits are reserved for sandboxing. In Cage’s prototype, we assign three bits for internal memory safety and one bit for enforcing external memory safety. This isolates a single WASM instance while assigning three bits for internal memory safety.
Lastly, Cage must ensure that adding an untrusted pointer to the heap base does not allow WASM code to craft arbitrary values to escape the sandbox. To provide these guarantees, Cage masks the WASM index before address computation, as shown in Fig. 13. Cage masks bits 56–59 when only external memory safety is enabled (Fig. 13(a)) and bit 56 when both internal and external memory safety are enabled (Fig. 13(b)).
Limitations. A natural limitation of Cage is the number of sandboxes within a single process. MTE provides a limited number of tags (16). Cage reserves the zero tag for the runtime, leaving the remaining 15 for the sandboxes. In a future version of Cage, the number of available sandboxes can be increased by ensuring tags are only reused for memory regions outside the address range reachable by WASM pointers, i.e., via combining guard pages with memory tagging.
We must exclude certain tags from being generated by MTE instructions, as Cage uses the tag bit 56 to distinguish between runtime and guest memory. On top of that, Cage must also exclude the zero tag from the guest, as this tag is reserved for guard slots and untagged segments. To this end, at runtime startup, we specify which tags can be generated using the prctl mechanism.
| Variant | Ptr width | Internal | External | Ptr auth |
| baseline wasm32 | 32-bit | No | No | No |
| baseline wasm64 | 64-bit | No | No | No |
| Cage-mem-safety | 64-bit | Yes | No | No |
| Cage-ptr-auth | 64-bit | No | No | Yes |
| Cage-sandboxing | 64-bit | No | Yes | No |
| Cage | 64-bit | Yes | Yes | Yes |
7. Evaluation
7.1. Experimental Setup
We conduct our benchmarks on a Google Pixel 8 equipped with an ArmV9 Tensor G3 chip, including one Cortex-X3 (2.91 GHz, out-of-order), four Cortex-A715 (2.37 GHz, out-of-order), and four Cortex-A510 (1.7 GHz, in-order) cores. All cores feature MTE and PAC with FEAT_FPAC enabled, making PAC trap on authentication.
Methodology. We evaluate Cage on the PolyBench/C 3.2 suite (Pouchet, [n. d.]) and microbenchmarks with the configurations outlined in Table 3. These include wasm32 and wasm64 as baselines and the memory-safety, pointer-authentication (6.3), and sandboxing (6.4) components from Cage, as well as all components of Cage combined. We perform and compare each benchmark on every type of CPU core available on the Tensor G3 by pinning benchmarks to a single core.
7.2. Performance Overheads
Figure 14 illustrates the mean runtime overheads for the PolyBench/C benchmarks.
Cage memory safety. Compared to wasm64, our memory safety extension has a mean overhead of , , and .
[…]
Cage. When combining our memory safety extension with MTE-based sandboxing, we see a mean speedup of , , and compared to wasm64, while providing stronger security guarantees.
Cage pointer authentication. As the PolyBench/C test suite does not exercise virtual function calls, we only see an overhead within error margins over 64-bit WASM. We perform a microbenchmark, comparing static against dynamic and authenticated dynamic function calls in Fig. 15. We observe virtually no overhead occurring as the result of pointer authentication, as adding pointer authentication only adds 5 cycles of latency, which is not noticeable. Switching from static to dynamic function calls results in a 15 %–22 % overhead, depending on the CPU being run on.
Cage Startup Overhead. We measure the startup overhead for a WASM instance with a static memory of 128 MiB, calling a function that immediately returns. The overhead of tagging the linear memory is hidden by the runtimes startup overhead.
7.3. Memory Overheads
The memory overheads of Cage consist of two factors: (i) the overhead of switching from 32-bit to 64-bit WASM and (ii) the overhead of memory tags caused by MTE, which is of the memory with MTE enabled. Tags are stored in a separate physical address space, the tag PA space (ARM Ltd., [n. d.]). This space is not managed by the operating system and is not included in maximum resident set size (rss) measurements. We thus estimate the memory overhead with MTE by factoring in 4 bits for every 16 byte of memory, which results in an additional . As our maximum rss measurements include the runtimes memory, which does not use MTE, this should be considered an over-approximation; in reality, the memory overhead of Cage is lower. We measure the mean overhead of wasm64 over wasm32 at and thus estimate Cage’s overhead at .
7.4. Security Guarantees
We evaluate the security guarantees from the perspective of internal and external memory safety (Section 4.1).
External memory safety. Cage implements sandboxing with MTE in synchronous mode, which enforces sandboxes through hardware. This prevents sandbox escapes, such as CVE2023-26489, from accessing memory outside the linear memory. We prevent malicious code from forging tags to escape the sandbox, as described in 6.4. However, we limit the number of sandboxes in one process to at most 15, which is required to assign a distinct tag to each sandbox and one to the runtime.
Internal memory safety.
Our choice to utilize a memory-tagging-based approach does not provide complete memory safety, as we rely on a limited number of tags.
As we reserve one tag for guard allocations and untagged segments, the probability for a tag collision is .
If we utilize MTE for sandboxing, the chance of a tag collision rises to .
Cage deterministically protects against off-by-one overflows, use-after-free, and double-free errors, which are caught at least until the reuse of a memory allocation.
Signing function pointers further reduces the possible attack surface.
As WASM already provides strong guarantees against control-flow-based attacks, adding pointer authentication primitives does not substantially improve security.
Cage protects against reusing leaked function pointers between instances and statically deducing function pointers.
We recommend Cage to be used as a secondary defense mechanism to mitigate classes described in Table 2.
With the ability to deploy the prototype in production, bugs may be found in production environments not discovered by testing workloads.
| Variant | Instr | Granule | Sets 0 | memset |
| memset | - | - | No | Yes |
| stg | stg | 16 bytes | No | No |
| st2g | st2g | 32 bytes | No | No |
| stgp | stgp | 16 bytes | Yes | No |
| stzg | stzg | 16 bytes | Yes | No |
| st2zg | st2zg | 32 bytes | Yes | No |
| stg+memset | stg | 16 bytes | Yes | Yes |
| st2g+memset | st2g | 32 bytes | Yes | Yes |
Initializing tagged memory. We benchmark several primitives to initialize an uncached 128 MiB large memory region while setting its allocation tag, representing the scenario of setting up a WASM instance. We run the configurations in Table 4 with synchronous MTE and measure the results in Fig. 16. stzg, stz2g, and stgp are slightly faster than a raw memset, even though they initialize memory and set tags. We believe this is the case as they do not perform tag checks before accessing memory (ARM Ltd., [n. d.]). The Cortex-X3 core compensates for its lower throughput (Table 1) with a higher clock speed (2.91 GHz) compared to the Cortex-A715 (2.37 GHz).
8. Related Work
Prior research exists on enriching WASM with memory safety properties. A significant project is MS-WASM (Disselkoen et al., 2019; Michael et al., 2023), a memory safety extension for WASM that introduces a new segment memory distinct from the linear memory, preventing access through arbitrary offsets. The segment memory relies on accesses through unforgeable handles, akin to CHERI pointers (Woodruff et al., 2014). On the contrary, Cage does not introduce a distinct memory region and a new pointer type, but rather builds upon 64-bit WASM without altering its memory layout to be able to minimize its runtime overheads on devices supporting Arm’s MTE.
RichWasm is another approach towards memory safety for WASM (Paraskevopoulou et al., 2024). It provides a richly typed intermediate language for safe memory interactions between languages with varying memory management models. RichWasm allows for static detection of memory safety violations in the context of mixed-language interoperability between strongly typed languages like Rust or OCaml. Unsafe languages (e.g., C/C++) lack information for static safety analysis and are not directly supported by RichWasm’s type-driven safety model. In contrast, Cage targets languages that do not encode safety in their runtime or type system.
Wasm memcheck (wmemcheck) (Alliance, 2024a) is a tool in wasmtime providing the ability to check for invalid mallocs, double-frees, reads, and writes inside a WASM module, assuming certain properties about malloc and free. It is conceptually similar to Valgrind memcheck (Nethercote and Seward, 2007).
9. Conclusion
In this paper, we present Cage, a hardware-accelerated WebAssembly toolchain to provide memory safety, consisting of a minimally invasive and adaptable WASM extension, a compiler toolchain based on LLVM, a modified wasi-libc that includes a custom allocator to provide spatial and temporal memory safety, and a wasmtime implementation, that is responsible for compiling and running the Cage’s WASM extension using MTE and PAC. Further, Cage improves the performance of WASM sandboxing mechanism by utilizing MTE as a replacement for software-based bounds checks. Our evaluation of Cage highlights that Cage is a memory safety solution for WASM that is suitable for production deployment as it incurs minimal runtime and memory overheads while providing efficient memory safety guarantees.
Software artifact. Cage will be publicly available with its entire experimental setup.
Acknowledgement
We thank the anonymous reviewers for their helpful feedback. We also thank Fritz Rehde, Janne Mantyla, and Carlos Chinea Perez for their work and feedback in the early stages of the project. This work was partially supported through a research grant from Huawei Research Finland, the TUM Innovation Network Resilient, Trustworthy, Sustainable (ReTruSt), and an ERC starting grant (ID: 101077577).
Appendix A Artifact Appendix
A.1. Abstract
This appendix provides the necessary information to build the artifacts and reproduce the experiments of the CGO’25 paper “Cage: Hardware-Accelerated Safe WebAssembly” by M. Fink, D. Stavrakakis, D. Sprokholt, S. Chakraborty, J.-E. Ekberg., and P. Bhatotia. Cage provides a memory safety abstraction for WebAssembly, with an implementation in LLVM to transparently compile unmodified C/C++ programs, a modified libc that provides memory safety for heap allocations, and an implementation in wasmtime that utilizes Arm MTE and PAC to implement the abstraction. We provide an artifact including pre-compiled binaries and all sources and scripts to build and evaluate the artifact to reproduce the results and figures in this paper.
A.2. Artifact Check-List (Meta-Information)
Program: LLVM with Cage modifications, source included; wasmtime with Cage modifications, source included; wasi-libc with Cage modifications, source included.
Compilation: Clang 17, rustc 1.80
Transformations: Stack allocation hardening as an LLVM pass.
Binary: Pre-built binaries for wasmtime, LLVM, wasi-sdk included. Source code and makefiles to re-generate binaries included.
Run-time environment: Provided binaries built for Linux 6.8.12 (nixOS) and Android 14.
Hardware: We require an AArch64 device with both PAC and MTE (Pixel 8) and an x86-64 machine for cross-compilation.
Metrics: Average runtime overhead; estimated memory overhead
Output: PDFs for the plots; Text files with raw data for tables and claims in text.
Experiments: Makefile to run all experiments is included.
How much disk space required (approximately)?: 25 GiB.
How much time is needed to prepare workflow (approximately)?: 2–3 hours.
How much time is needed to complete experiments (approximately)?: 2–3 days.
Publicly available?: Yes.
Code licenses (if publicly available)?: Apache License with LLVM Exceptions (LLVM, wasi-libc); Apache License (wasmtime).
Workflow framework used?: Makefiles.
Archived (provide DOI)?: 10.5281/zenodo.13347823
A.3. Description
A.3.1. How delivered
All source code can be found at the git repositories below, as well as the following persisted DOI for the artifact: 10.5281/zenodo.13347823, which contains all source code as well as the scripts to build and evaluate all artifacts.
A.3.2. Hardware Dependencies
The evaluation is performed on a Google Pixel 8 with Arm MTE and PAC. We cross-compile LLVM, wasmtime, and the benchmarks on an x86 machine.
A.3.3. Software Dependencies
We require the following set of software on the Google Pixel 8 device:
Termux
sshd
bash
We require the following set of software on the x86 machine to compile LLVM, wasmtime, and evaluate the benchmarks:
Linux (tested with 6.8.12)
nix (tested with 2.18.5): All other dependencies are fetched and pinned to a specific version using the nix package manager and can be found in the nix/ directory in the artifact.
A.3.4. Benchmarks
We run the PolyBench/C benchmark suite to measure the runtime and memory overhead of Cage’s components compared to the 32- and 64-bit baselines. To measure startup overheads, we measure the overhead of instantiating a module declaring a 128 MiB memory and calling an empty function. To measure pointer authentication overheads, we measure a modified version of PolyBench/C’s 2mm benchmark, where the matrix multiplication is moved into a function call that is either performed statically or dynamically through a vtable.
A.4. Installation
To get started, download the artifact from Zenodo, navigate to the artifact directory, and run the following command to download all required dependencies using nix.
A.5. Evaluation and Expected Result
Expected duration: [1 human-minute + 2–3 compute-days]
Running the experiments on the Pixel 8 devices takes a long time. This is primarily caused by our choice to run all experiments on all three types of cores found in the Pixels chipset. Running the benchmarks on the low-power Cortex-A510 cores takes up most of the runtime. To perform the evaluation, run: